# Towards verifications of Krylov complexity

### v1

@@ Abstract
Krylov complexity is considered to provide a measure of the growth of operators evolving under Hamiltonian dynamics. The main strategy is the analysis of the structure of Krylov subspace $\mathcal{K}_{M}(\mathcal{H},\eta)$ spanned by the multiple applications of the Liouville operator $\mathcal{L}$ defined by the commutator in terms of a Hamiltonian $\mathcal{H}$, $\mathcal{L}:=[\mathcal{H},\cdot]$ acting on an operator $\eta$, $\mathcal{K}_{M}(\mathcal{H},\eta)=\text{span}\{\eta,\mathcal{L}\eta,\ldots,\mathcal{L}^{M-1}\eta\}$. For a given inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) of the operators, the orthonormal basis $\{\mathcal{O}_{n}\}$ is constructed from $\mathcal{O}_{0}=\eta/\sqrt{(\eta,\eta)}$ by Lanczos algorithm. The moments $\mu_{m}=(\mathcal{O}_{0},\mathcal{L}^{m}\mathcal{O}_{0})$ are closely related to the important data $\{b_{n}\}$ called Lanczos coefficients. I present the exact and explicit expressions of the moments $\{\mu_{m}\}$ for 16 quantum mechanical systems which are exactly solvable both in the Schrödinger and Heisenberg pictures. The operator $\eta$ is the sinusoidal coordinate which is the variable of the eigenpolynomials $\{P_{n}(\eta)\}$ of the exactly solvable Hamiltonian $\mathcal{H}$.
@@ 1 Introduction
In the study of quantum chaos, Krylov complexity is proposed as a measure of the growth of operators in the Heisenberg picture [1, 2, 3, 4, 5, 6]. The definition of Krylov complexity is universal, that is, applicable to a very complicated system as well as an extremely ordered system, for example, an exactly solvable system in the Heisenberg picture. As materials to support the validity of the concept of Krylov complexity by contrast [7, 8], I present one of the most basic ingredients of the theory, the moments $\{\mu_{m}\}$ (2.20) of a special operator $\eta$ called sinusoidal coordinate of many exactly solvable quantum systems in the Heisenberg picture. They are ten discrete quantum mechanical systems of finite dimensions and two infinite ones plus four ordinary one dimensional quantum mechanics. In all these systems, the sinusoidal coordinate is the variable in the eigenpolynomials $\{P_{n}(\eta)\}$ of the Hamiltonian $\mathcal{H}$. These polynomials all belong to the Askey scheme of hypergeometric orthogonal polynomials [9, 10, 11].
This paper is organised as follows. In section 2, the basic concepts of Krylov complexity are briefly recapitulated [1] through the orthonormalisation of operators in a Krylov subspace [12] by Lanczos algorithm [13]. In section 3 the outline of exactly solvable discrete quantum systems in the Schrödinger picture is briefly reviewed [14]. Corresponding solutions of Heisenberg equation of motion [15, 16], which are the principal tool of the present paper, are discussed in some detail in section 4. The main results of the paper, the exact and explicit expressions of the moments $\{\mu_{m}\}$ are derived for ten exactly solvable discrete quantum mechanics in section 5 and four ordinary one-dimensional quantum mechanics in section 6. The inner product including the Boltzmann factor is introduced in section 6 to deal with unbounded Hamiltonians. The exact expressions of the moments for the bounded and unbounded Hamiltonians look very similar as shown in two main Theorems 5.2 and 6.1. The final section is for a summary and some comments.
@@ 2 Orthonormalisation of operators in the Krylov subspace
in which $\dagger$ denotes the hermitian conjugation and *** means the complex conjugation. Throughout this paper the operator $\eta$ is assumed to be a real diagonal matrix,
$\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\text{diag}\{\eta(0),\eta(1),\ldots,\eta(N)\},\quad\eta(x)\in\mathbb{R},\ x\in\mathcal{X}.$ (2.5)
@@
$(\mathcal{V},\mathcal{W})\stackrel{{\scriptstyle\text{def}}}{{=}}\text{Tr}[\![\mathcal{V}^{\dagger}\mathcal{W}]\!]=\sum_{x\in\mathcal{X}}(\!(x|\mathcal{V}^{\dagger}\mathcal{W}|x)\!)=\sum_{n\in\mathcal{X}}\langle n|\mathcal{V}^{\dagger}\mathcal{W}|n\rangle=(\mathcal{W},\mathcal{V})^{*},$ (2.8)
@@
$\mathcal{W}_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{0},\quad b_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{0}|\!|=|\!|\mathcal{L}\mathcal{O}_{0}|\!|,\quad\mathcal{O}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{1}}\mathcal{W}_{0}.$ (2.13)
By the above property (2.11), $\mathcal{O}_{1}$ is automatically orthogonal with the previous $\mathcal{O}_{0}$,
@@
$\mathcal{W}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{1}-b_{1}\mathcal{O}_{0},\quad b_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{1}|\!|,\quad\mathcal{O}_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{2}}\mathcal{W}_{1}.$ (2.15)
@@
$\mathcal{W}_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{k}-b_{k}\mathcal{O}_{k-1},\quad b_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{k}|\!|,\quad\mathcal{O}_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{k+1}}\mathcal{W}_{k}.$ (2.17)
@@
The orthonormalisation is complete when all the Lanczos coefficients $\{b_{n}\}$ are determined. The orthonormalisation stops at $\mathcal{O}_{k}$ when $b_{k+1}$ vanishes, $b_{k+1}=0$. Two explicit examples of the stopped orthonormalisation will be shown in §5.1. Since the Hilbert space is $\mathbb{C}^{N+1}$, the totality of the basis $\{O_{n}\}$ is less than $(N+1)^{2}$.
@@
[…]
@@
In some work, a formula involving the determinant of Hankel matrix of moments
@@
In a monograph [17] Viswanath and Müller reported a recursive formula to determine $\{b_{1}^{2},\ldots,b_{K}^{2}\}$ based on the knowledge of $\mu_{2},\ldots,\mu_{2K}$. The formula was recapitulated in [1](A.5).
@@
[…]
@@
[…]
@@
In the rest of this paper I present the explicit forms of the Heisenberg operator solution (2.24) and the moments $\{\mu_{2m}\}$ of more than dozen quantum mechanical systems. Based on the exact knowledge of the moments, the Lanczos coefficients $\{b_{n}\}$ and the functions $\{\varphi_{n}(t)\}$ (2.23) can be evaluated as precisely as wanted.
@@ 3 Exactly solvable discrete quantum mechanics
$\begin{split}&\text{(i)}\ \eta(x)=x,\ K,H,\quad\text{(ii)}\ \eta(x)=x(x+d),\ dH,R\\ &\text{(iii)}\ \eta(x)=1-q^{x},\qquad\text{(iv)}\ \eta(x)=q^{-x}-1,\ qH,qqK,qK,aqK\\ &\text{(v)}\ \eta(x)=(q^{-x}-1)(1-dq^{x}),\quad dqH,qR\end{split}$ (3.5)
@@
$\begin{split}&\text{(i)}\ \mathcal{E}(n)=n,\ K,dH,\quad\text{(ii)}\ \mathcal{E}(n)=n(n+d),\ H,R\\ &\text{(iii)}\ \mathcal{E}(n)=1-q^{n},\ qqK\qquad\text{(iv)}\ \mathcal{E}(n)=q^{-n}-1,\ dqH,aqK\\ &\text{(v)}\ \mathcal{E}(n)=(q^{-n}-1)(1-dq^{n}),\quad qK,qH,qR.\end{split}$ (3.6)
@@ 4 Solutions of Heisenberg equation of motion
As shown in (2.24), the explicit form of the Heisenberg operator solution is essential for the determination of the functions $\{\varphi_{n}(t)\}$. Although the Heisenberg solution of the harmonic oscillator potential ($x^{2}$) was known in the early days of quantum mechanics, it was late 1970’s that those for four other potentials in one-dimensional quantum mechanics were reported by Nieto and Simmons [19]. They were for the potentials $x^{2}+1/x^{2}$, $1/\sin^{2}x$, $-1/\cosh^{2}x$ and the Morse potential. The term ‘sinusoidal coordinate’ was also coined by them, meaning that $\eta$ undergoes sinusoidal motion with frequencies depending on the energy. About a quarter century after Nieto and Simmons, the list of exact Heisenberg operator solutions was enlarged by Odake and myself to include many discrete quantum mechanics [14, 15, 16, 20] and some multi-particle dynamics [21].
The essence is the discovery [16] (this paper will be cites as I) that the Hamiltonian $\mathcal{H}$ and the sinusoidal coordinate $\eta$ of exactly sovable systems in the Schrödinger picture all satisfy a simple commutation relation
@@
in which $R_{0}(\mathcal{H})$, $R_{1}(\mathcal{H})$ and $R_{-1}(\mathcal{H})$ are polynomials in $\mathcal{H}$ of maximal degree 2, 1 and 2, respectively, reflecting the power counting of $\mathcal{H}$ on both sides. Those $R_{i}(\mathcal{H})$’s may contain some system parameters but not dynamical operators other than $\mathcal{H}$. In terms of the Liouville operator $\mathcal{L}$, the obove commutation relation reads
$\mathcal{L}^{2}\eta=\eta\,R_{0}(\mathcal{H})+\mathcal{L}\eta\,R_{1}(\mathcal{H})+R_{-1}(\mathcal{H}).$ (4.2)
@@
$\mathcal{L}^{m}\eta=\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\quad m\in\mathbb{N}_{\geq 0},$
@@
$\mathcal{L}^{m}\eta=$ $\,\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\qquad m\in\mathbb{N}_{\geq 0},$ (4.4)
@@
$\hskip 113.81102pt\Longrightarrow\langle n|\eta|n\rangle=-\frac{R_{-1}\bigl{(}\mathcal{E}(n)\bigr{)}}{R_{0}\bigl{(}\mathcal{E}(n)\bigr{)}},\qquad n\in\mathcal{X}.$ (4.15)

### v2

@@ Abstract
Krylov complexity is considered to provide a measure of the growth of operators evolving under Hamiltonian dynamics. The main strategy is the analysis of the structure of Krylov subspace $\mathcal{K}_{M}(\mathcal{H},\eta)$ spanned by the multiple applications of the Liouville operator $\mathcal{L}$ defined by the commutator in terms of a Hamiltonian $\mathcal{H}$, $\mathcal{L}:=[\mathcal{H},\cdot]$ acting on an operator $\eta$, $\mathcal{K}_{M}(\mathcal{H},\eta)=\text{span}\{\eta,\mathcal{L}\eta,\ldots,\mathcal{L}^{M-1}\eta\}$. For a given inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) of the operators, the orthonormal basis $\{\mathcal{O}_{n}\}$ is constructed from $\mathcal{O}_{0}=\eta/\sqrt{(\eta,\eta)}$ by Lanczos algorithm. The moments $\mu_{m}=(\mathcal{O}_{0},\mathcal{L}^{m}\mathcal{O}_{0})$ are closely related to the important data $\{b_{n}\}$ called Lanczos coefficients. I present the exact and explicit expressions of the moments $\{\mu_{m}\}$ for 16 quantum mechanical systems which are exactly solvable both in the Schrödinger and Heisenberg pictures. The operator $\eta$ is the variable of the eigenpolynomials. Among them six systems show a clear sign of ‘non-complexity’ as vanishing higher Lanczos coefficients $b_{m}=0$, $m\geq 3$.
Published: Progress of Theoretical and Experimental Physics,
Volume 2024, Issue 6, June 2024, 063A01,
Doi: https://doi.org/10.1093/ptep/ptae073
@@ 1 Introduction
In the study of quantum chaos, Krylov complexity is proposed as a measure of the growth of operators in the Heisenberg picture [1, 2, 3, 4, 5, 6, 7]. The definition of Krylov complexity is universal, that is, applicable to a very complicated system as well as an extremely ordered system, for example, an exactly solvable system in the Heisenberg picture. As materials to support the validity of the concept of Krylov complexity by contrast [8, 9, 10], I present one of the most basic ingredients of the theory, the moments $\{\mu_{m}\}$ (2.20) of a special operator $\eta$ called sinusoidal coordinate of many exactly solvable quantum systems in the Heisenberg picture. They are ten discrete quantum mechanical systems of finite dimensions and two infinite ones plus four ordinary one dimensional quantum mechanics. In all these systems, the eigenvalues of the Hamiltonian are simple explicit functions of the system parameters and the sinusoidal coordinate is the variable in the eigenpolynomials $\{P_{n}(\eta)\}$ of the Hamiltonian $\mathcal{H}$. These polynomials all belong to the Askey scheme of hypergeometric orthogonal polynomials [11, 12, 13].
This paper is organised as follows. In section 2, the basic concepts of Krylov complexity are briefly recapitulated [1] through the orthonormalisation of operators in a Krylov subspace [14] by Lanczos algorithm [15]. In section 3 the outline of exactly solvable discrete quantum systems in the Schrödinger picture is briefly reviewed [16]. Corresponding solutions of Heisenberg equation of motion [17, 18], which are the principal tool of the present paper, are discussed in some detail in section 4. The main results of the paper, the exact and explicit expressions of the moments $\{\mu_{m}\}$ are derived for ten exactly solvable discrete quantum mechanics in section 5 and four ordinary one-dimensional quantum mechanics in section 6. The inner product including the Boltzmann factor is introduced in section 6 to deal with unbounded Hamiltonians. The exact expressions of the moments for the bounded and unbounded Hamiltonians look very similar as shown in two main Theorems 5.2 and 6.1. Among these 16 exactly solvable systems, six systems related to the Krawtchouk and dual Hahn §5.1, Meixner and Charlier §6.1, Hermite and Laguerre §6.2 polynomials show a very clear sign of ‘non-complexity’ as vanishing higher Lanczos coefficients $b_{m}=0$, $m\geq 3$. These six systems all share a linear spectrum $\mathcal{E}(n)\propto n$. The final section is for a summary and some comments.
@@ 2 Orthonormalisation of operators in the Krylov subspace
in which $\dagger$ denotes the hermitian conjugation and * means the complex conjugation. Throughout this paper the operator $\eta$ is assumed to be a real diagonal matrix,
$\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\text{diag}\{\eta(0),\eta(1),\ldots,\eta(N)\},\quad\eta(x)\in\mathbb{R},\quad x\in\mathcal{X}.$ (2.5)
@@
$(\mathcal{V},\mathcal{W})\stackrel{{\scriptstyle\text{def}}}{{=}}\text{Tr}[\![\mathcal{V}^{\dagger}\mathcal{W}]\!]=(\mathcal{W}^{\dagger},\mathcal{V}^{\dagger})=\sum_{x\in\mathcal{X}}(\!(x|\mathcal{V}^{\dagger}\mathcal{W}|x)\!)=\sum_{n\in\mathcal{X}}\langle n|\mathcal{V}^{\dagger}\mathcal{W}|n\rangle=(\mathcal{W},\mathcal{V})^{*},$ (2.8)
@@
$\mathcal{W}_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{0},\quad b_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{0}|\!|=|\!|\mathcal{L}\mathcal{O}_{0}|\!|,\quad\text{if}\ \ b_{1}\neq 0\ \Rightarrow\ \mathcal{O}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{1}}\mathcal{W}_{0},$ (2.13)
and it goes on until a zero norm operator is produced. By the above property (2.11), $\mathcal{O}_{1}$ is automatically orthogonal with the previous $\mathcal{O}_{0}$,
@@
$\mathcal{W}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{1}-b_{1}\mathcal{O}_{0},\quad b_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{1}|\!|,\quad\text{if}\ \ b_{2}\neq 0\ \Rightarrow\ \mathcal{O}_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{2}}\mathcal{W}_{1}.$ (2.15)
@@
$\mathcal{W}_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{k}-b_{k}\mathcal{O}_{k-1},\quad b_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{k}|\!|,\quad\text{if}\ \ b_{k+1}\neq 0\ \Rightarrow\ \mathcal{O}_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{k+1}}\mathcal{W}_{k}.$ (2.17)
@@
The orthonormalisation is complete when all the Lanczos coefficients $\{b_{n}\}$ are determined. The orthonormalisation stops at $\mathcal{O}_{k}$ when $b_{k+1}$ vanishes, $b_{k+1}=0$. Two explicit examples of the stopped orthonormalisation will be shown in §5.1. Since the Hilbert space is $\mathbb{C}^{N+1}$, the totality of the basis $\{\mathcal{O}_{n}\}$ is less than $(N+1)^{2}$.
@@
$\mu_{2m}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{O}_{0},\mathcal{L}^{2m}\mathcal{O}_{0}),\quad 0\leq m\leq n,\quad\mu_{0}=1.$ (2.20)
@@
In some work [1](A.4), a formula involving the determinant of Hankel matrix of moments
@@
In a monograph [19] Viswanath and Müller reported a recursive formula to determine $\{b_{1}^{2},\ldots,b_{K}^{2}\}$ based on the knowledge of $\mu_{2},\ldots,\mu_{2K}$ [19](3.33) and its reverse, from $\mu_{2},\ldots,\mu_{2K}$ to $\{b_{1}^{2},\ldots,b_{K}^{2}\}$, [19](3.34). These formulas were recapitulated in [1](A.5),(A.7).
@@
$K(\mathcal{H},\eta;t)\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{n}n\varphi_{n}(t)^{2},\quad\varphi_{n}(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\bigl{(}i^{n}\mathcal{O}_{n},\mathcal{O}(t)\bigr{)}\in\mathbb{R},\quad n=1,2,\ldots,$ (2.23)
@@
$\mathcal{O}(t)\stackrel{{\scriptstyle\text{def}}}{{=}}e^{i\mathcal{H}t}\mathcal{O}_{0}e^{-i\mathcal{H}t}=e^{i\mathcal{H}t}\eta e^{-i\mathcal{H}t}/|\!|\eta|\!|,\quad\bigl{(}\mathcal{O}(t),\mathcal{O}(t)\bigr{)}=1.$ (2.24)
@@
In the rest of this paper I present the explicit forms of the Heisenberg operator solution (2.24) and the moments $\{\mu_{2m}\}$ of more than a dozen quantum mechanical systems. Based on the exact knowledge of the moments, the Lanczos coefficients $\{b_{n}\}$ and the functions $\{\varphi_{n}(t)\}$ (2.23) can be evaluated as precisely as wanted.
@@ 3 Exactly solvable discrete quantum mechanics
$\begin{split}&\text{(i)}\ \eta(x)=x,\ K,H,\quad\text{(ii)}\ \eta(x)=x(x+d),\ dH,R,\\ &\text{(iii)}\ \eta(x)=1-q^{x},\qquad\text{(iv)}\ \eta(x)=q^{-x}-1,\ qH,qqK,qK,aqK,\\ &\text{(v)}\ \eta(x)=(q^{-x}-1)(1-dq^{x}),\quad dqH,qR.\end{split}$ (3.5)
@@
$\begin{split}&\text{(i)}\ \mathcal{E}(n)=n,\ K,dH,\quad\text{(ii)}\ \mathcal{E}(n)=n(n+d),\ H,R,\\ &\text{(iii)}\ \mathcal{E}(n)=1-q^{n},\ qqK,\qquad\text{(iv)}\ \mathcal{E}(n)=q^{-n}-1,\ dqH,aqK,\\ &\text{(v)}\ \mathcal{E}(n)=(q^{-n}-1)(1-dq^{n}),\quad qK,qH,qR.\end{split}$ (3.6)
@@ 4 Solutions of Heisenberg equation of motion
As shown in (2.24), the explicit form of the Heisenberg operator solution is essential for the determination of the functions $\{\varphi_{n}(t)\}$. Although the Heisenberg solution of the harmonic oscillator potential ($x^{2}$) was known in the early days of quantum mechanics, it was late 1970’s that those for four other potentials in one-dimensional quantum mechanics were reported by Nieto and Simmons [21][24]. They were for the potentials $x^{2}+1/x^{2}$, $1/\sin^{2}x$, $-1/\cosh^{2}x$ and the Morse potential. The term ‘sinusoidal coordinate’ was also coined by them, meaning that $\eta$ undergoes sinusoidal motion with frequencies depending on the energy. About a quarter century after Nieto and Simmons, the list of exact Heisenberg operator solutions was enlarged by Odake and myself to include many discrete quantum mechanics [16, 17, 18, 25] and some multi-particle dynamics [26].
The essence is the discovery [18] (this paper will be cited as I) that the Hamiltonian $\mathcal{H}$ and the sinusoidal coordinate $\eta$ of exactly sovable systems in the Schrödinger picture all satisfy a simple commutation relation
@@
in which $R_{0}(\mathcal{H})$, $R_{1}(\mathcal{H})$ and $R_{-1}(\mathcal{H})$ are polynomials in $\mathcal{H}$ of maximal degree 2, 1 and 2, respectively, reflecting the power counting of $\mathcal{H}$ on both sides. Those $R_{i}(\mathcal{H})$’s may contain some system parameters but not dynamical operators other than $\mathcal{H}$. In terms of the Liouville operator $\mathcal{L}$, the above commutation relation reads
$\mathcal{L}^{2}\eta=\eta\,R_{0}(\mathcal{H})+\mathcal{L}\eta\,R_{1}(\mathcal{H})+R_{-1}(\mathcal{H}),$ (4.2)
@@
$\mathcal{L}^{m}\eta=\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\quad m\in\mathbb{Z}_{\geq 0},$
@@
$\mathcal{L}^{m}\eta=$ $\,\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\qquad m\in\mathbb{Z}_{\geq 0},$ (4.4)
@@
$\hskip 113.81102pt\Longrightarrow\langle n|\eta|n\rangle=-\frac{R_{-1}\bigl{(}\mathcal{E}(n)\bigr{)}}{R_{0}\bigl{(}\mathcal{E}(n)\bigr{)}},\qquad n\in\mathcal{X},$ (4.15)